How many sheets of a dielectric are necessary to make a condenser having a capacity of 5.5nF when the sheets are 15cm x 30cm in area and 0.29cm thick? Use k = 8



1
Expert's answer
2022-04-11T17:04:01-0400

The capacitance of the parallel plate capacitor

"C=\frac{\kappa\epsilon_0A}{d}"

Hence, the thickness of dielectric

"d=\frac{\kappa\epsilon_0A}{C}"

"d=\frac{8*8.85*10^{-12}*0.15*0.30}{5.5*10^{-9}}=5.8*10^{-4}\:\rm m"

The number of sheets

"N=\frac{d}{d_1}=\frac{5.8*10^{-4}\:\rm m}{29*10^{-4}\:\rm m}=0.2"